"The sum of two numbers is 90. The larger number is 14 more than 3 times the smaller number. Find the numbers."
Answer by TimothyLamb(4379)

i + j = 90
j = 3i + 14
---
i + j = 90
i + 3i + 14 = 90
4i = 76
---
i = 19
j = 71
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php